Sahara is3 times as old as rob. Their ages together add up to 16. How old is sahara? Show me how to write out the equation so i can get the answer.
Answer by nerdybill(7384)

Sahara is3 times as old as rob. Their ages together add up to 16. How old is sahara? Show me how to write out the equation so i can get the answer.
.
Let s = Sahara's age
and r = Rob's age
.
From:"Sahara is3 times as old as rob." we get equation 1:
s = 3r
.
From:"Their ages together add up to 16" we get equation 2:
s+r = 16
.
Substitute 's' as defined in equation 1 into equation 2 and solve for r:
s+r = 16
3r+r = 16
4r = 16
r = 4 years old (Rob's age)
.
Sahara's age:
s = 3r = 3(4) = 12 years old